OpenStudy (anonymous):

How do you simplify the following expression: x^-2y^-5/ -4x^4y^-3

OpenStudy (dumbcow):

\[\frac{x^{-2}y^{-5}}{-4x^{4}y^{-3}} = -\frac{1}{4x^{4+2}y^{-3+5}} = -\frac{1}{4x^{6}y^{2}}\]

OpenStudy (dumbcow):

its not showing the neg signs in the exponents for some reason

OpenStudy (anonymous):

thank you sooooooo much, I thought that was how it was done, but for some reason I didn't think I did it right

OpenStudy (dumbcow):

your welcome